# A Class of De Giorgi Type and Hölder Continuity for Some Problems in Musielak-Orlicz-Sobolev Spaces

### v1

@@ Abstract.
In this paper, we introduce a new class of De Giorgi type functions, denoted by $\mathcal{B}_{G(x,t)}$, and establish the Hölder continuity of its elements under suitable additional assumptions on the generalized N-function $G(x,t)$. As an application, we prove the Hölder continuity of solutions to quasilinear equations whose principal part is in divergence form with $G(x,t)$-growth conditions, including both critical and standard growth cases. The novelty of our work lies in the generalization of the Hölder continuity results previously known for variable exponent [30, X, Fan and D. Zhao] and Orlicz [51, G. M. Lieberman] problems. Moreover, our results encompass a wide variety of quasilinear equations.
@@ 1.1. Overview
$\operatorname{div}A(x,u,\nabla u)=B(x,u,\nabla u)\quad\text{in }\Omega.$ (1.1)
@@
$\Delta_{p}u:=\operatorname{div}(|\nabla u|^{p-2}\nabla u).$
@@
These ideas and techniques were systematically extended to tackle the more general non-polynomial growth of operators in the Orlicz setting. The foundational work by Donaldson and Trudinger [25] first established the boundedness ($L^{\infty}$) of solutions. Lieberman [51], who proved that weak solutions to quasilinear equations of the form $\operatorname{div}(a(|\nabla u|)\nabla u)=0$ are locally Hölder continuous ($C^{0,\alpha}$). This required the associated N-function $G(t)$ to satisfy the condition $\Delta_{2}$, which is essential for the Sobolev-Poincaré inequalities and the scaling arguments of the De Giorgi-Nash-Moser theory. These contributions laid the foundation of the regularity theory in the Orlicz framework.
Although Orlicz spaces encompass many classical examples, there remain important cases that they do not cover. For this reason, considerable research has focused on the variable exponent framework, where the growth is governed by a function $p(x)$. The analysis of problems involving terms of the type $t^{p(x)}$, initiated by Fan and Zhao [30], revealed a new regularity phenomenon. They proved that weak solutions of equations such as
$\operatorname{div}(|\nabla u|^{p(x)-2}\nabla u)=0,$ are locally Hölder continuous, provided the exponent function $p(x)$ is log-Hölder continuous. This condition, expressed as
@@
was shown to be not only sufficient but essentially necessary for regularity, a fact later solidified by the work of Acerbi and Mingione [3].
@@
$\operatorname{div}\!\left(|\nabla u|^{p-2}\nabla u+a(x)|\nabla u|^{q-2}\nabla u\right),$
@@
Due to the vast scope and rapid development of regularity theory, we cannot cite all contributions individually. We acknowledge with gratitude the many valuable works not mentioned here and instead refer the reader to comprehensive surveys and bibliographies, such as [57, 58]. For detailed discussions of specific regularity results, we direct the reader to the following foundational works and the extensive literature contained therein: for the classical Laplacian and linear theory, see [21, 73, 70, 71]; for the $p$-Laplacian and $(p,q)$-growth, see [39, 38, 8, 37, 80, 45, 48, 50, 79, 9, 10, 46, 19, 47, 52, 53, 56, 54, 1, 2, 32, 36, 33, 27, 42, 78]; for Orlicz and other non-standard growth conditions, see [51, 40, 18, 26, 12, 11, 59]; for double-phase problems, see [15, 16, 7, 20]; and for variable exponent $p(x)$-Laplacian problems, see [30, 31, 3, 17, 43].
The Musielak-Orlicz growth framework represents the most general setting, encompassing all the previously mentioned cases: standard $p$-growth, Orlicz, variable exponent, and double-phase (or $(p,q)$) growth. To the best of our knowledge, the literature contains only one other work specifically devoted to the Hölder continuity of solutions to divergence-form problems under such general Musielak-Orlicz growth conditions, namely that of Wang, Liu, and Zhao [81]. In their analysis, the authors impose a key structural assumption on the function $G(x,t)$:
There exists a strictly increasing differentiable function $\xi:[0,+\infty)\to[0,+\infty)$ that satisfies
@@
A significant limitation of this condition is its restricted applicability. The assumption (A11), in particular, is not satisfied by many fundamental models central to the theory of non-standard growth problems. Notably, it fails to hold for variable exponent spaces, $(p,q)$-growth equations, and double-phase functionals, precisely the paradigmatic cases that a comprehensive Musielak-Orlicz theory should seek to unify. Consequently, while providing an interesting result under a specific set of constraints, the framework of [81] does not offer the generality required to recover these critical classical and modern examples.
@@ 1.2. Assumptions and Main Results
For every $x_{0}\in\Omega$ and all sufficiently small $R>0$, there exists $L_{0}>1$ such that for all $t>1$:
@@
where $r(x,t):=\dfrac{tg(x,t)}{G(x,t)}$.
@@
The constants $a_{i}$, $b_{j}$, and $c_{k}$ are all positive, and we fix parameters $F_{1}>1$ and $F_{2}>1$. We also introduce the generalized N-functions:
@@
$g^{+}\leq h_{i}^{-}\leq\frac{h_{i}(x,t)\,t}{H_{i}(x,t)}\leq h_{i}^{+},$ $\qquad\text{for all }x\in\Omega,\ t>0,\quad i=1,2,$ ($\mathcal{H}_{2}$)
@@
$\max_{x\in\Omega}|u(x)|\leq M.$ ($\mathcal{M}$)
@@ 1.3. Examples
defines a generalized N-function that satisfies the structural assumptions (G0)–(G3); see [41, Section 7.1]. Therefore, our results also apply to the variable exponent setting, which has been extensively studied by X. Fan and D. Zhao [30].
@@ 2.2. Musielak-Orlicz-Sobolev space $W^{1,G(x,t)}$
    For a.a. $x\in\Omega$, the function $G(x,t)$ is even, continuous, increasing, and convex in $t$, and for each $t\in\mathbb{R}$, $G(x,t)$ is measurable in $x$;
@@
    $\lim_{t\to\infty}\frac{G(x,t)}{t}=\infty$, for a.a. $x\in\Omega$;
@@
$G(x,2t)\leq C_{0}G(x,t)+\varphi(x),\quad\text{for a.a. }x\in\Omega\text{ and all }t\geq 0.$
@@
    $\mu\leq G^{-1}(x,1)\leq\frac{1}{\mu},\quad\text{for a.a. }x\in\mathbb{R}^{n};$
@@
    for every $t\in[1,1/|B|]$, for a.a. $x,y\in B$, and every ball $B\subset\mathbb{R}^{n}$ with $|B|\leq 1$;
@@
    $G_{1}(x,t)\leq C_{1}G_{2}(x,C_{2}t)+\varphi(x),\quad\text{for a.a. }x\in\Omega\text{ and all }t\geq 0.$
@@
For any generalized N-function $G(x,t)$, the function $\widetilde{G}:\Omega\times\mathbb{R}\to\mathbb{R}$ defined by
$\widetilde{G}(x,t):=\sup_{s\geq 0}\left(ts-G(x,s)\right),\quad\text{for all }x\in\Omega\text{ and }t\geq 0,$ (2.2)
is called the complementary function of $G$.
@@
Let $G(x,t)$ be a generalized N-function. Suppose that $t\mapsto g(x,t)$ is continuous and increasing on $\mathbb{R}$ for a.a. $x\in\Omega$. Moreover, assume that there exist constants $g^{-},g^{+}\in\mathbb{R}$ such that
@@
    $\min\{s^{g^{-}},s^{g^{+}}\}G(x,t)\leq G(x,st)\leq\max\{s^{g^{-}},s^{g^{+}}\}G(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
@@
    $\min\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t)\leq\widetilde{G}(x,st)\leq\max\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
@@
The subsequent proposition deals with some topological properties of the Musielak-Orlicz space, see [72, Theorem 7.7 and Theorem 8.5].
@@
$W^{1,G(x,t)}(\Omega):=\left\{u\in L^{G(x,t)}(\Omega):\ |\nabla u|\in L^{G(x,t)}(\Omega)\right\}.$
@@
$\|u\|:=\|u\|_{L^{G(x,t)}(\Omega)}+\|\nabla u\|_{L^{G(x,t)}(\Omega)},\ \ \text{for all}\ u\in W^{1,G(x,t)}(\Omega),$ (2.9)
where $\|\nabla u\|_{L^{G(x,t)}(\Omega)}:=\||\nabla u|\|_{L^{G(x,t)}(\Omega)}$.
@@
$G^{\ast}(x,t)=G(x,N^{-1}(x,t)),\text{ for a.a. }x\in\mathbb{R}^{n}\text{ and all }t\geq 0,$
@@
    $\min\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t)\leq G^{*}(x,st)\leq\max\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t),\text{ for a.a. }x\in\Omega$  $\text{ and all }s,\ t\geq 0$;

### v2

@@ Abstract.
In this paper, we introduce a new class of De Giorgi type functions, denoted by $\mathcal{B}_{G(x,t)}$, and establish the Hölder continuity of its elements under suitable additional assumptions on the generalized N-function $G(x,t)$. As an application, we prove the Hölder continuity of solutions to quasilinear equations whose principal part is in divergence form with $G(x,t)$-growth conditions, including both critical and standard growth cases. The novelty of our work lies in the generalization of the Hölder continuity results previously known for variable exponent [27, X, Fan and D. Zhao] and Orlicz [48, G. M. Lieberman] problems. Moreover, our results encompass a wide variety of quasilinear equations.
@@ 1.1. Overview
$\operatorname{div}A(x,u,Du)=B(x,u,Du)\quad\text{in }\Omega.$ (1.1)
@@
$\Delta_{p}u:=\operatorname{div}(|Du|^{p-2}Du).$
@@
These ideas and techniques were systematically extended to tackle the more general non-polynomial growth of operators in the Orlicz setting. The foundational work by Donaldson and Trudinger [22] first established the boundedness ($L^{\infty}$) of solutions. Lieberman [48], who proved that weak solutions to quasilinear equations of the form $\operatorname{div}(a(|Du|)Du)=0$ are locally Hölder continuous ($C^{0,\alpha}$). This required the associated N-function $G(t)$ to satisfy the condition $\Delta_{2}$, which is essential for the Sobolev-Poincaré inequalities and the scaling arguments of the De Giorgi-Nash-Moser theory. These contributions laid the foundation of the regularity theory in the Orlicz framework.
Although Orlicz spaces encompass many classical examples, there remain important cases that they do not cover. For this reason, considerable research has focused on the variable exponent framework, where the growth is governed by a function $p(x)$. The analysis of problems involving terms of the type $t^{p(x)}$ was initiated by Zhikov; see [79, 80]. Subsequently, Fan and Zhao [27] uncovered a new regularity phenomenon by showing that weak solutions of equations of the form
$\operatorname{div}\big(|Du|^{p(x)-2}Du\big)=0$
are locally Hölder continuous, provided that the exponent function $p(x)$ is log-Hölder continuous. This condition, namely
@@
was shown to be not only sufficient but essentially necessary for regularity, a fact later clarified and reinforced by the work of Acerbi and Mingione [3].
@@
$\operatorname{div}\!\left(|Du|^{p-2}Du+a(x)|Du|^{q-2}Du\right),$
@@
Due to the vast scope and rapid development of regularity theory, we cannot cite all contributions individually. We acknowledge with gratitude the many valuable works not mentioned here and instead refer the reader to comprehensive surveys and bibliographies, such as [54, 55]. For detailed discussions of specific regularity results, we direct the reader to the following foundational works and the extensive literature contained therein: for the classical Laplacian and linear theory, see [18, 70, 67, 68]; for the $p$-Laplacian and $(p,q)$-growth, see [34, 77, 45, 47, 76, 43, 16, 44, 49, 50, 53, 51, 1, 2, 29, 33, 30, 24, 40, 75]; for Orlicz and other non-standard growth conditions, see [48, 11, 38, 37, 39, 35, 15, 23, 9, 8, 56]; for double-phase problems, see [12, 13, 7, 17, 11]; and for variable exponent $p(x)$-Laplacian problems, see [27, 28, 3, 14, 41].
The Musielak–Orlicz growth framework represents the most general setting, as it encompasses all previously studied models, including the standard $p$-growth, Orlicz growth, variable exponent growth, and double-phase (or $(p,q)$) growth. A major advance in the regularity theory for generalized Orlicz growth was achieved by Harjulehto, Hästö, and Lee [39], who established Hölder continuity of local minimizers of integral functionals with Musielak–Orlicz growth under very general structural assumptions, namely (A0), (A1), and (A2) (see Subsection 2.2). In particular, their results do not require any logarithmic continuity condition on the growth function, such as (G3) (see Subsection 1.2), and they cover important models including the double-phase case. This highlights the strength and generality of the variational approach developed in [39] for the regularity of minimizers, which relies on comparison estimates and harmonic approximation techniques.
The present paper is concerned with a different class of problems. We study weak solutions of quasilinear elliptic equations and systems in divergence form (1.1), possibly including lower-order terms, for which a variational structure is not available in general. Consequently, the methods developed in [39] cannot be applied directly in our setting. Instead, we adopt a De Giorgi–type iteration adapted to Musielak–Orlicz growth conditions. Within this non-variational framework, an additional control on the spatial oscillation of the effective growth, formulated in assumption (G3), is required in order to derive the key structural estimates and to obtain Hölder regularity of solutions. We also emphasize that our analysis covers Neumann boundary conditions, which are not addressed in the variational framework of [39].
To the best of our knowledge, the literature contains only one other work specifically devoted to the Hölder continuity of weak solutions to divergence-form elliptic problems under Musielak–Orlicz growth conditions, namely the paper by Wang, Liu, and Zhao [78]. In their analysis, the authors impose a restrictive structural assumption on the growth function $G(x,t)$. More precisely, they assume the existence of a strictly increasing differentiable function $\xi:[0,+\infty)\to[0,+\infty)$ satisfying
@@
A significant limitation of this framework is that assumption (A11) excludes several fundamental models in the theory of non-standard growth. In particular, it is not satisfied by variable exponent growth, $(p,q)$-growth equations, or double-phase functionals, which are among the main motivating examples of the Musielak–Orlicz setting. As a consequence, while the results of [78] are obtained under a specific and interesting set of hypotheses, their approach does not provide a unified treatment of these central growth models.
@@ 1.2. Assumptions and Main Results
There exists $L_{0}>1$ such that, for every $x_{0}\in\Omega$, for all sufficiently small $R>0$ and for all $t>1$, we heve
@@
where $r(x,t):=\dfrac{tg(x,t)}{G(x,t)}$. For the definition of $G^{-1}$ see Definition 2.3.
@@
For the definition of $\widetilde{H_{1}}^{-1}$ see Definitions 2.3 and 2.7. The constants $a_{i}$, $b_{j}$, and $c_{k}$ are all positive, and we fix parameters $F_{1}>1$ and $F_{2}>1$. We also introduce the generalized N-functions:
@@
$g^{+}\leq h_{i}^{-}\leq\frac{h_{i}(x,t)\,t}{H_{i}(x,t)}\leq h_{i}^{+}\leq g_{*}^{-}:=\frac{ng^{-}}{n-g^{-}},$ $\qquad\text{for all }x\in\Omega,\ t>0,\quad i=1,2,$ ($\mathcal{H}_{2}$)
@@
$\mathop{\mathrm{ess\,sup}}_{x\in\Omega}|u(x)|\leq M.$ ($\mathcal{M}$)
@@ 1.3. Examples
defines a generalized N-function that satisfies the structural assumptions (G0)–(G3); see [36, Section 7.1]. Therefore, our results also apply to the variable exponent setting, which has been extensively studied by X. Fan and D. Zhao [27].
@@
The following remark clarifies the logical relationship between assumptions (G2) and (G3). Although both conditions impose continuity-type requirements on the generalized $\mathrm{N}$-function $G(x,t)$ with respect to the spatial variable, they control fundamentally different aspects of the growth. Condition (G2) is a local comparability assumption, which is mainly used to derive functional-analytic tools such as Poincaré and Sobolev-type inequalities in Musielak–Orlicz spaces. In contrast, condition (G3) provides a quantitative control on the spatial oscillation of the effective growth exponent $\frac{tg(x,t)}{G(x,t)}$, which plays a crucial role in De Giorgi–type iteration arguments. The remark below shows, by means of explicit examples, that these two assumptions are independent in general.
Remark 1.15.
Conditions (G2) and (G3) are independent in general.
(1) (G2) does not imply (G3). Consider the variable exponent case
$G(x,t)=t^{p(x)},$
where $p:\Omega\to\mathbb{R}$ is bounded and measurable but not log-Hölder continuous. Then condition (G2) holds (see, e.g., [36]), since $G^{-1}(x,t)=t^{1/p(x)}$ is locally comparable on balls. However,
$\frac{tg(x,t)}{G(x,t)}=p(x),$
and thus condition (G3) fails whenever $p(\cdot)$ lacks log-Hölder continuity. Hence (G2) does not imply (G3).
(2) (G3) does not imply (G2). Let
$G(x,t)=t^{p}(1+a(x)),\quad p>1,$
where $a:\Omega\to[0,\infty)$ is unbounded or highly oscillatory. Then
$\frac{tg(x,t)}{G(x,t)}=p,$
so condition (G3) holds trivially. However, the lack of local comparability of $G(x,t)$ prevents (G2) from being satisfied in general.
Therefore, (G2) and (G3) are logically independent assumptions.
@@ 2.2. Musielak-Orlicz-Sobolev space $W^{1,G(x,t)}$
    For a.e. $x\in\Omega$, the function $G(x,t)$ is even, continuous, strictly increasing, and convex in $t$, and for each $t\in\mathbb{R}$, $G(x,t)$ is measurable in $x$;
@@
    $\lim_{t\to\infty}\frac{G(x,t)}{t}=\infty$, for a.e. $x\in\Omega$;
@@
Let $G:\Omega\times[0,\infty)\to[0,\infty)$ be a generalized $\mathrm{N}$-function. Since for a.e. $x\in\Omega$ the mapping $t\mapsto G(x,t)$ is continuous and strictly increasing on $0,\infty)$, we can define $G^{-1}(x,\cdot)$ as the continuous inverse of $G(x,\cdot)$. In particular, it holds that
$G^{-1}(x,G(x,t))=G(x,G^{-1}(x,t))=t,\quad\text{for a.e.\ }x\in\Omega\text{ and all }t\geq 0.$
Definition 2.4.
@@
$G(x,2t)\leq C_{0}G(x,t)+\varphi(x),\quad\text{for a.e. }x\in\Omega\text{ and all }t\geq 0.$
@@
    $\mu\leq G^{-1}(x,1)\leq\frac{1}{\mu},\quad\text{for a.e. }x\in\mathbb{R}^{n};$
@@
    for every $t\in[1,1/|B|]$, for a.e. $x,y\in B$, and every ball $B\subset\mathbb{R}^{n}$ with $|B|\leq 1$;
@@
    $G_{1}(x,t)\leq C_{1}G_{2}(x,C_{2}t)+\varphi(x),\quad\text{for a.e. }x\in\Omega\text{ and all }t\geq 0.$
@@
Let $G:\Omega\times[0,\infty)\to[0,\infty)$ be a generalized $\mathrm{N}$-function. The function $\widetilde{G}:\Omega\times[0,\infty)\to[0,\infty)$ defined by
$\widetilde{G}(x,t):=\sup_{s\geq 0}\bigl\{ts-G(x,s)\bigr\},\quad\text{for a.e.\ }x\in\Omega\text{ and all }t\geq 0,$ (2.2)
is called the complementary function (or conjugate function) of $G$. Moreover, $\widetilde{G}$ is also a generalized $\mathrm{N}$-function.
@@
Let $G(x,t)$ be a generalized N-function. Suppose that $t\mapsto g(x,t)$ is continuous and increasing on $\mathbb{R}$ for a.e. $x\in\Omega$. Moreover, assume that there exist constants $g^{-},g^{+}\in\mathbb{R}$ such that
@@
    $\min\{s^{g^{-}},s^{g^{+}}\}G(x,t)\leq G(x,st)\leq\max\{s^{g^{-}},s^{g^{+}}\}G(x,t),\text{ for a.e. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
@@
    $\min\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t)\leq\widetilde{G}(x,st)\leq\max\{s^{\widetilde{g^{-}}},s^{\widetilde{g^{+}}}\}\widetilde{G}(x,t),\text{ for a.e. }x\in\Omega$  $\text{ and all }s,\ t\geq 0.$
@@
The subsequent proposition deals with some topological properties of the Musielak-Orlicz space, see [69, Theorem 7.7 and Theorem 8.5].
@@
$W^{1,G(x,t)}(\Omega):=\left\{u\in L^{G(x,t)}(\Omega):\ |Du|\in L^{G(x,t)}(\Omega)\right\}.$
@@
$\|u\|:=\|u\|_{L^{G(x,t)}(\Omega)}+\|Du\|_{L^{G(x,t)}(\Omega)},\ \ \text{for all}\ u\in W^{1,G(x,t)}(\Omega),$ (2.9)
where $\|Du\|_{L^{G(x,t)}(\Omega)}:=\||Du|\|_{L^{G(x,t)}(\Omega)}$.
@@
$G^{\ast}(x,t)=G(x,N^{-1}(x,t)),\text{ for a.e. }x\in\mathbb{R}^{n}\text{ and all }t\geq 0,$
@@
    $\min\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t)\leq G^{*}(x,st)\leq\max\left\{s^{g^{-}_{*}},s^{g^{+}_{*}}\right\}G^{*}(x,t),\text{ for a.e. }x\in\Omega$  $\text{ and all }s,\ t\geq 0$;